# Augmenting Smart Contract Decompiler Output through Fine-grained Dependency Analysis and LLM-facilitated Semantic Recovery

### v1

@@ 4.3. Correctness Verification
Program-behavior Equivalence Check.  SmartHalo aims to check the program-behavior equivalence between methods of original decompiled code and methods of optimized decompiled code. We use $m$ and $m^{\prime}$ to represent the two versions of the same method in the original decompiled code and optimized decompiled code.
@@
SmartHalo conducts the symbolic execution on methods $m$ and $m^{\prime}$ for generating symbolic summary $s$ and $s^{\prime}$. If methods $m$ and $m^{\prime}$ are equivalent in terms of program behaviors, their symbolic summary $s$ and $s^{\prime}$ are also equivalent in terms of functionality. An equivalence assertion is a first-order logic formula $\Phi$ that is used to assert such equivalence (Person et al., 2008).
(1) $\Phi=\neg(s\Leftrightarrow s^{{}^{\prime}})$
@@
SmartHalo utilize formal verification method to analyze symbolic summaries $s$ and $s^{\prime}$, and determine whether equivalence assertion $\Phi$ is satisfied. If $\Phi$ is satisfied, method $m$ is non-equivalent to method $m^{\prime}$ in terms of program behaviors, otherwise, they are equivalent to each other in terms of program behaviors. Further, SmartHalo integrates the formal verification method that relies on SMT solver (i.e., Z3 solver (de Moura and Bjørner, 2008)), to find out satisfying assignments for equivalence assertion $\Phi$.
@@
Note that since the symbolic summaries generally contain non-linear constraints, it is difficult for the Z3 solver to solve them. Such non-linear constraints commonly result in the solving failure. We eliminate most of the solving failure by utilizing methods proposed in ARDiff (Badihi et al., 2020).
@@
Static Rule Violation Check.  SmartHalo traverse throughout the optimized code, then leverages violation rejection rules to identify and reject incorrect variable types predicted by the LLM.
@@
Here, we summarize the violation rejection rules  [^2] More details about the violation rejection rules can be found in our project repository integrated by SmartHalo. Each violation rejection rule is composed of two parts, including specific premises (contents above the line) and conclusions (contents below the line). They are organized as follows.
(2) $\pi\vdash e:\theta$
@@
Finally, the incorrect optimization results predicted by LLM would be integrated with verification suggestions to form the violation information. SmartHalo retransmits the violation information into LLM to infer the correct method boundaries, variable types, and contract attributes anew. Further, the interaction can be iterated continually between LLM-driven semantic enrichment and correctness verification, until SmartHalo amends all the optimization errors identified by correctness verification, or a maximum iteration limit is reached.
5. Evaluation
@@
    Can the output of SmartHalo help reverse engineers understand the smart contract bytecode in the real world?
5.1. Implementation and Evaluation Setup
@@
Implementation.  We utilize ChatGPT (GPT-3.5) to support the LLM component of SmartHalo, due to the significantly expensive overhead of GPT fee. Additionally, we also conduct a small-scale experiments ( Section 5.2) to evaluate the effectiveness of SmartHalo that integrates the versions of GPT-4o, which shows SmartHalo indeed performs better with more advanced LLM models. Note that SmartHalo is designed to be model-agnostic and unlimited to a certain large language model, it offers the flexibility to integrate alternative large language models.
SmartHalo was implemented with around 1,799 lines of code in Python 3.8.10. All evaluation experiments were conducted on an Ubuntu 20.04 server equipped with an Intel i9-10980XE CPU (3.0GHz), an RTX 3090 GPU, and 250GB of RAM.
Evaluation Dataset and Ground-truth Establishment.  Previous research (Chen et al., 2022a; Lacomis et al., 2019) take significantly more effort to construct large datasets that encompass the different kinds of source code, as they are heavily dependent on the diversity of the training data. In contrast, one of the advantages of SmartHalo is its integration with LLMs (e.g., ChatGPT), which enables it to achieve strong performance without the need for extensive model training. Hence, we only require a relatively small testing dataset, rather than a large training dataset, to evaluate the effectiveness of SmartHalo.
@@
Our evaluation dataset derives from a largest open-source smart contract dataset  (xbl, [n. d.]) as of Apirl 2024, totaling 963,151 smart contracts. From this original dataset, we randomly sampled 500 methods to construct the evaluation dataset. Ultimately, we filter out the methods the decompiler cannot handle, and obtain 456 pairs of source code and decompiler outputs of functions, for forming the final evaluation dataset. Here, we utilize method pairs instead of entire contracts because GPT integrated by SmartHalo requires shorter code snippets as input.
@@
Furthermore, we established the ground truth for the evaluation dataset. Specifically, the ground truth for method boundaries and variable types was determined by manually comparing the decompiled code with the source code, as these aspects are explicitly stated in the source code. In contrast, the ground truth for contract attributes was established through manual inspection by our domain experts (i.e., researchers with three years of experience in smart contracts), who annotated the contract attributes via majority voting. As mentioned earlier, contract attributes refer to the semantic meaning represented by state variables (e.g., asset, address). To ensure comprehensive coverage of contract attribute categories, we also randomly sampled an additional 1,000 smart contracts from the original dataset and used a Large Language Model (LLM) to analyze the usage patterns of state variables within them. With the usage patterns, our experts manually summarize the representative categories as [ Limit, Fee, Flag, Address, Asset, Router, Others ].
@@
  • Method boundary match. We confirm the method boundary prediction is correct, by manually investigating whether the predicted method boundaries totally match the ground truth method boundaries.
  • Type match. Similar to DIRTY(Chen et al., 2022a), We confirm the type prediction is correct, if and only if the predicted types are totally matched the ground truth type. We serialize types to strings and determine the type matching by the string matching.
  • Contract attribute precision. We determine the precision of contract attributes by determining whether the predicted attributes are fully consistent with the ground truth attributes.
  • Recompilation failures. This metric represents errors yielded by the compiler when we are recompiling decompiled outputs. This indicates bugs or limitation in decompiler output (Williams-King et al., 2020).
5.2. Overall Effectiveness
To address RQ1, we evaluate the overall effectiveness of SmartHalo by comparing the optimized code generated by SmartHalo with the original decompiler output produced by the Dedaub decompiler (i.e., the web version of Gigahorse (Grech et al., 2019)) using four key metrics. To achieve this, we run SmartHalo on the evaluation dataset. Additionally, we conduct a detailed analysis of the reasons behind optimization failures. As can be seen in Table 2, the optimized code generated by SmartHalo significantly improves all four metrics compared to the original decompiler output. For example, the precision of method boundary identification is enhanced by 30.03%, and the precision of variable type inference is increased by 12.01%. Note that the the total number of variable types is inconsistent between Dedaub decompiler and SmartHalo, because the optimized code correctly recovers more methods which generate more parameter types. Moreover, while the original decompiler output lacks any contract attribute information, SmartHalo successfully recovers 68.06% of contract attributes. Additionally, unlike the original decompiler output, which is not suitable for recompilation, 60.22% of the optimized code generated by SmartHalo can be recompiled using a compiler.
Table 2. Comparison results between the optimized code generated by SmartHalo and original decompiler output produced by the Dedaub decompiler, for evaluate the overall effecctiveness of SmartHalo.
Method border
Variable type
Contract attribute
@@ 5.2. Overall Effectiveness
Original decompiler output (Baseline)
50.48% (316/626)
62.25% (676/1086)
N/A 100%
SmartHalo
80.51% (504/626)
84.26% (1349/1601)
68.06% (752/1102)
39.78% (185/465)
“NA” : the corresponding tool does not support optimizing this deficiency.
Table 3. Comparsion results between SmartHalo with GPT-3.5 and SmartHalo with GPT-4o.
Method border
Variable type
Contract attribute
SmartHalo with GPT-3.5
78.99% (94/119)
71.52% (113/158)
73.13% (98/134)
SmartHalo with GPT-4o
87.39% (104/119)
90.39% (254/281)
80.65% (100/124)
Optimization failures.  We manually inspect all the optimization failures of the four metrics. We found that most of optimization failures are caused by the limited model capability of GPT-3.5. As proved by our experiment in Table 3, SmartHalo equipped with GPT-4o can achieve precision rates of 87.39% for method boundaries, 90.39% for variable types, and 80.65% for contract attributes.
Exploration on GPT-4o.  Here, we explore how the model capability of LLM contributes to the overall effectiveness of SmartHalo. The study of the kind of LLM that can better drive SmartHalo is not the focus of our work. Conversely, we aim to explore how a more advanced LLM can further enhance the effectiveness of SmartHalo. GPT-4o is a more advanced LLM and performs better in various tasks compared with GPT-3.5. Hence, we compare the overall effectiveness in terms of three metrics between SmartHalo equipped with GPT-3.5 and SmartHalo equipped with GPT-4o. To save the expensive GPT fees, we randomly sample 50 method pairs from the evaluation dataset. We run the SmartHalo equipped with GPT-4o and SmartHalo equipped with GPT-3.5 overall the 50 method pairs. As can be seen in Table 3, SmartHalo equipped with GPT-4o further enhances the overall effectiveness. In conclusion, SmartHalo can integrate LLMs with better model capability to further enhance its effectiveness.
5.3. Comparison to Prior Work
To address RQ2, we compare SmartHalo with recent works. It’s important to note that we do not conduct comparative experiments regarding method boundary recovery, as there is a lack of existing work on optimizing method boundaries for smart contracts. In addition, SigRec (Chen et al., 2022b) is not open-sourced, and the model dataset for DeepInfer is unavailable for download (despite our inquiries with the authors, we received no response). Fortunately, we contacted the authors of SmartDagger (Liao et al., 2022) and obtained the artifact. Hence, we run SmartHalo and SmartDagger over the evaluation dataset to compare their precision in terms of contract attribute.
As can be seen in Table 4, SmartHalo significantly outperforms SmartDagger. For example, the precision of the contract attribute is improved by 44.69%. Further, we manually investigate the optimization failures generated by SmartDagger (Liao et al., 2022). Our findings indicate that the performance of SmartHalo degrades when encountering our dataset, which includes many newly-emerging smart contracts, due to limitations in its model capability and training dataset. Conversely, SmartHalo equipped with GPT-3.5 can maintain a higher precision of 68.06% for contract attributes. When equipped with GPT-4o, SmartHalo achieves even higher precision, with values of 80.65%.
@@ 5.3. Comparison to Prior Work
Table 4. Comparison results between SmartHalo and SmartDagger in terms of contract attribute
Contract attribute
SmartDagger 10.07% (111/1102)
SmartHalo 68.06% (752/1102)
5.4. Ablation Study

### v2

@@ IV-C Correctness Verification
Program-behavior Equivalence Check.  SmartHalo aims to check the program-behavior equivalence between functions of original decompiled code and functions of optimized decompiled code. We use $m$ and $m^{\prime}$ to represent the two versions of the same function in the original decompiled code and optimized decompiled code.
@@
SmartHalo conducts the symbolic execution on functions $m$ and $m^{\prime}$ for generating symbolic summary $s$ and $s^{\prime}$. If functions $m$ and $m^{\prime}$ are equivalent in terms of program behaviors, their symbolic summary $s$ and $s^{\prime}$ are also equivalent in terms of functionality. An equivalence assertion is a first-order logic formula $\Phi$ that is used to assert such equivalence [33].
$\Phi=\neg(s\Leftrightarrow s^{{}^{\prime}})$ (1)
@@
SmartHalo utilize formal verification method to analyze symbolic summaries $s$ and $s^{\prime}$, and determine whether equivalence assertion $\Phi$ is satisfied. If $\Phi$ is satisfied, function $m$ is non-equivalent to function $m^{\prime}$ in terms of program behaviors, otherwise, they are equivalent to each other. Further, SmartHalo integrates the formal verification method that relies on SMT solver (i.e., Z3 solver [34]), to find out satisfying assignments for equivalence assertion $\Phi$.
@@
Note that since the symbolic summaries generally contain non-linear constraints, it is difficult for the Z3 solver to solve them. Such non-linear constraints commonly result in the solving failure. We eliminate most of the solving failure by utilizing methods proposed in ARDiff [35].
@@
Rule-based Type Violation Check.  SmartHalo traverse throughout the optimized code, then leverages violation rejection rules to identify and reject incorrect variable types predicted by the LLM.
@@
Here, we summarize the violation rejection rules integrated by SmartHalo in Figure 7. Each violation rejection rule is composed of two parts, including specific premises (contents above the line) and conclusions (contents below the line). They are organized as follows.
$\pi\vdash e:\theta$ (2)
@@
Finally, the incorrect optimization results predicted by LLM would be integrated with verification suggestions to form the violation information. SmartHalo retransmits the violation information into LLM to infer the correct function boundaries, variable types, and contract attributes anew. Further, the interaction can be iterated continually between LLM-driven semantic enrichment and correctness verification, until SmartHalo amends all the errors identified by correctness verification, or a maximum iteration limit is reached.
V Evaluation
@@
    Can SmartHalo generalize to real-world complex contracts, different decompilers or different LLMs?
  5. RQ5.
    What is the efficiency of SmartHalo in terms of both runtime and monetary costs?
  6. RQ6.
    How effective is the output of SmartHalo in terms of helping downstream program analysis task?
V-A Implementation and Evaluation Setup
@@
Implementation.  SmartHalo was implemented with around 1,799 lines of code in Python 3.8.10. We utilize ChatGPT (GPT-3.5) to support the LLM component of SmartHalo, due to the significantly expensive overhead of GPT fee. Note that SmartHalo is designed to be model-agnostic and unlimited to a certain large language model, it offers the flexibility to integrate alternative large language models. All evaluation experiments were conducted on an Ubuntu 20.04 server equipped with an Intel i9-10980XE CPU (3.0GHz), an RTX 3090 GPU, and 250GB of RAM.
@@
Evaluation Dataset and Ground-truth Establishment.  Our evaluation dataset derives from a largest open-source smart contract dataset [17] as of Apirl 2024, totaling 963,151 smart contracts. From this original dataset, we randomly sampled 500 functions to construct the evaluation dataset. Ultimately, we filter out the functions the decompiler cannot handle, and obtain 456 pairs of source code and decompiler outputs of functions, for forming the final evaluation dataset. Noting that evaluation dataset size is similar to those used in SOTA studies [4, 18]. Here, we utilize function pairs instead of entire contracts because GPT integrated by SmartHalo requires shorter code snippets as input.
Furthermore, we established the ground truth for the evaluation dataset. Specifically, the ground truth for function boundaries and variable types was determined by directly comparing the decompiled code with the source code, as these aspects are explicitly stated in the source code. In contrast, the ground truth for contract attributes was established through manual summarization by our domain researchers. To mitigate bias, the manual analysis procedure was conducted in a rigorous manner. The process consists of three steps. In the first step, six researchers with at least two years of relevant domain experience were invited and organized into three pairs. Among these, two pairs were designated as annotators, while the remaining pair, composed of researchers with at least four years of experience, served as referee experts. In the second step, each annotator within a pair independently conducts the manual investigation. The third step is cross-validation and quality control. For each pair of annotators, they perform the cross-validation and discuss the labels until they obtain a consensus on the classification results. If consensus cannot be reached, the referee experts are consulted to provide the final decision. In addition, we calculate the average Cohen’s Kappa coefficient to measure the consistency of labeling results, which is 0.772, thus indicating a substantial agreement between each pair of researchers.
@@
As mentioned earlier, contract attributes refer to the semantic meaning represented by state variables (e.g., asset, address). To ensure comprehensive coverage of contract attribute categories, we also randomly sampled an additional 1,000 smart contracts from the original dataset and used a Large Language Model (LLM) to analyze the usage patterns of state variables within them. With the usage patterns, our experts manually summarize the representative categories as [Limit, Fee, Flag, Address, Asset, Router, Others].
TABLE II: Distribution and characteristic of the evaluation dataset.
Metric Evaluation dataset
Average lines 163
Functionalities
Cross-Chain interoperability, Market Mechanisms & Price Discovery,
Governance, Risk Management & Liquidation, Security & Consensus,
Token Management, Value Transfer & Incentive Distribution,
Entertainment & Probabilistic Gamebling
Solidity From 0.4.22 to 0.8.25
Deployment From April 2018 to April 2024
@@
  • function boundary match. We confirm the boundary prediction is correct for a given function, by manually investigating whether the starting and ending points are totally match those of sourcecode-level function.
  • Type match. Similar to DIRTY[16], We consider a type prediction to be correct only if the predicted type fully matches the ground truth type, including data layout, and the type and name of any fields if applicable.
  • Contract attribute match. Similar to type match, we determine a attribute prediction to be correct only if the predicted attribute fully matches the ground truth attribute.
  • Recompilation failures. It represents errors yielded by the compiler when recompiling decompiled outputs. This indicates bugs or limitation in decompiler output [36].
TABLE III: Comparison results between the optimized code generated by SmartHalo and original decompiler output produced by the Dedaub decompiler over the evaluation dataset, for evaluating the overall effecctiveness of SmartHalo.
Metrics Function boundary Variable type Contract attribute
@@
TP FP FN Prec. Recall TP FP FN Prec. Recall TP FP FN Prec. Recall Rate
Original
decompiler
316 149 310 67.96% 50.48% 676 410 975 62.25% 42.22% N/A 100%
SmartHalo 504 67 122 88.26% 80.51% 1349 113 252 92.27% 84.26% 752 350 75 68.06% 90.93% 39.78%
“N/A” : the corresponding tool does not support optimizing this deficiency.
V-B Overall Effectiveness
To address RQ1, we evaluate the overall effectiveness of SmartHalo by comparing the optimized code generated by SmartHalo with the original decompiler output produced by the Dedaub decompiler (i.e., the web version of Gigahorse [9]) using four key metrics. To achieve this, we run SmartHalo on the evaluation dataset. Additionally, we conduct a detailed analysis of the reasons behind optimization failures. As can be seen in Table III, the optimized code generated by SmartHalo significantly improves all four metrics compared to the original decompiler output. For example, the precision and recall of function boundary identification are enhanced by 20.30% and 30.03%, and the precision and recall of variable type inference are increased by 30.02% and 42.04%. Moreover, while the original decompiler output lacks any contract attribute information, SmartHalo successfully recovers the contract attributes with a precision of 68.06% and a recall of 90.93%. Additionally, unlike the original decompiler output, which is not suitable for recompilation, 60.22% of the optimized code generated by SmartHalo can be recompiled using a compiler.
Optimization failures.  We manually inspect all the optimization failures (i.e., false positives and false negatives) of the four metrics. Firstly, we leverage the open card sorting approach similar to prior research for the error taxonomy construction. We found that SmartHalo struggles when encountering the following types of functions or contracts: E1. low-level and delegated calls, E2. inline assembly, E3. off-chain reliance, E4. inheritance structure or E5. others. Then, we compute the rates for each type of optimization failure in the above-stated taxonomy. Lastly, we conduct the root-cause mapping between the SmartHalo’s optimization failure and core components (i.e., static analysis and LLM). We found that most of optimization failures in E1 are caused by the incorrect facts generated by Gigahorse (i.e., the decompiler of SmartHalo’s static analysis module). These optimization failures can be further eliminated by integrating a more accurate decompiler in the future. We found that most of optimization failures in E2 and E5 are caused by the limited model capability of GPT-3.5. As proved by our experiment in Table VII, SmartHalo equipped with GPT-4o mini can achieve a precision rate of 91.32% and a recall rate of 87.38% for function boundaries, a precision rate of 90.40% and a recall rate of 88.82% for variable types, and a precision rate of 80.66% and a recall rate of 91.78% for contract attributes. Any static analysis approach like ours is fundamentally limited to addressing optimization failures in E3 and E4, as optimization failures in E3 require analyzing the off-chain data source and optimization failures in E5 require analyzing the inheritance relationship which is actually missed at the bytecode level in smart contract.
@@ V-B Overall Effectiveness
TABLE IV: Comparison results between SmartHalo, VarLifter and SmartDagger in terms of contract attribute or variable type.
Metrics Contract attribute Variable type
TP FP FN Prec. Recall TP FP FN Prec. Recall
SmartDagger [5] 111 364 641 23.37% 10.07% N/A
VarLifter [6] N/A 115 31 1486 78.76% 7.18%
SmartHalo 752 350 75 68.06% 90.93% 1349 113 252 92.27% 84.26%
“N/A” means the corresponding tool does not support this optimization metric.
TABLE V: Comparison results between SmartHalo and LLM-facilitated approach.
Metrics Function boundary Variable type
TP FP FN Prec. Recall TP FP FN Prec. Recall
Original decompiler output (Baseline) 316 149 310 67.96% 50.48% 676 410 925 62.25% 42.22%
LLM without static inference prompt 321 144 305 69.03% 51.28% 375 144 1226 72.26% 23.42%
SmartHalo 504 67 122 88.26% 80.51% 1349 113 252 92.27% 84.26%
V-C Comparison to Prior Work
To address RQ2, we compare SmartHalo with recent works.
SigRec is not open-sourced, and we are unable to reproduce SigRec because its core algorithm (i.e., TASE) lacks disclosure of essential implementation details. The model dataset for DeepInfer is unavailable for download; moreover, the method omits critical information regarding model inputs, architecture, hyperparameters, and training protocols. Consequently, we could not reproduce DeepInfer. Finally, we were unable to conduct a comparative evaluation between SmartHalo and Neural‑FEBI, as the publicly released Neural‑FEBI repository omits key modules and contains erroneous model artifacts.
Indeed, SmartHalo is expected to outperforms SigRec, DeepInfer and Neural‑FEBI, as they can recover only a subset of variable types/function boundaries. Specifically, Neural‑FEBI does not support boundary recovery for complex functions such as modifier functions or functions inherited from other contracts/subcontracts. Conversely, SmartHalo optimizes boundaries for all functions. And SigRec and DeepInfer only partially recover known function signatures and parameter types. Conversely, SmartHalo optimizes types for all variables
Fortunately, we contacted the authors of SmartDagger [5] and obtained the artifact. Further, we download the artifact of VarLifter [6] from the open-source repository. Hence, we run SmartHalo and SmartDagger over the evaluation dataset to compare their precision and recall in terms of contract attribute. And we run SmartHalo and Varlifter over the evaluation dataset to compare their precision and recall in terms of variable type.
As can be seen in Table IV, SmartHalo significantly outperforms SmartDagger and Varlifter. In terms of contract attribute, compared to SmartDagger, the precision and recall of SmartHalo are improved by 44.69% and 80.86%. Further, we manually investigate the optimization failures generated by SmartDagger. Our findings indicate that the performance of SmartDagger degrades when encountering our dataset, which includes many newly-emerging smart contracts, due to limitations in its model capability and training dataset. In terms of variable type, compared to VarLifter, the precision and recall of SmartHalo are improved by 13.51% and 77.08%. Similarly, we also manually inspect the optimization failures generated by VarLifter. Our investigation results shows that VarLifter can only recover a small subset of variable types, owing to its low-coverage heuristic rules. And for most variable types, the output is marked as unknown, resulting in a large number of false negatives.
Conversely, SmartHalo equipped with GPT-3.5 achieves a higher precision of 68.06% and a higher recall of 90.93% for contract attributes, and exhibits a higher precision of 92.27% and a higher recall of 84.26% for variable types. When equipped with GPT-4o mini, SmartHalo achieves a even higher precision of 80.65% for contract attributes, and presents a more higher recall of 88.81% for variable types.
V-D Ablation Study
Actually, SmartHalo is composed of two components including static analysis and large language models. To answer RQ3, we compare the Dedaub decompiler, LLM (GPT-3.5) and SmartHalo in terms of function boundaries and variable types, to evaluate the contribution of two components to the overall effectiveness of SmartHalo. Note that we do not compare them in terms of contract attributes because LLM is incapable of outputting the contract attribute without our inference prompt (e.g., contract attribute candidate). To conduct this comparison, we further ran the LLM over the evaluation dataset.